Find the LCD then convert each rational expression into an equivalent rational expression with the LCD as the denominator.
5/12, 3/20

Answer by checkley77(12844)

5/12, 3/20
Factors of 12 are: 2*2*3
Factors of 20 are: 2*2*5
Thus the LCD=2*2*3*5=60
(5*5)/60, (3*3)/60
25/60, 9/60